A ball has a velocity vector V with components of Vx=2.40 m/s and Vy= - 1.70 m/s.
(a) What is the speed of the ball?
(b)When is the ball at 10.0 m on the X axis?
(c) When the ball has position of 10.0 meters on the X-axis, what is the position on the Y-axis?
1
Expert's answer
2020-11-25T10:51:04-0500

Answer(a)

"Speed\\v=\sqrt{(2.40)^2+(1.70)^2}\\v=2.94m/s"

(b)

"s_x=v_xt\\t=\frac{10}{2.40}=4.16s"

(c)

"s_y=v_yt\\s_y=1.70\times4.16\\s_y=7.07m"